Write the equation in slope-intercept form of the line that has a slope of 6 and contains the point (1, 1)
Answer by nyc_function(2645)

We use the point-slope formula.
y - y1 = m(x - x1)
y - 1 = 6(x - 1)
y - 1 = 6x - 6
y = 6x - 6 + 1
y = 6x - 5
==============================
By the way, the point-slope formula can be better understood as
y - y coordinate of your point = slope(x - x coordinate of your point)